The perimeter of a rectangular -shaped field is 138 ft. The width is 29 ft shorter than the length. Find the length and width of the field.
Let x =the length
Let y = the width
Write two equations. First complete the sentence that describes the formula for the perimeter of a rectangle.
Perimeter of a rectangle equals twice the length plus twice the width
Use the above formula and the assigned variables, x and y to write an equation.
2x+2y=138.
Recall the condition that the width of the rectangle is 29 ft shorter than the length. Now, write an equation for the width in terms of the length
y=

Answer by addingup(3677)

The perimeter of a rectangle is:
2L+2W= P The problem says the Width= L-29 and the Perimeter= 138:
2L+L-29+L-29= 138
4L-58= 138
4L= 196
L= 196/4
L= 49 This is the length, and the width:
W= 49-29= 20
Proof:
2(49)+2(20)= 138
98+40= 138
138= 138 We have the correct answer